# Linear response for skew-product maps with contracting fibres

### v1

@@ 1. Introduction
More precisely, given a family of dynamical systems $T_{\alpha}:M\to M$, with $\alpha$ taking values in an interval $J\subset\mathbb{R}$, it is natural to ask whether dynamical quantities are differentiable, or not, with respect to the parameter $\alpha$ in a suitable topology. In particular, if for each $\alpha$ the system $T_{\alpha}$ admits a unique physical measure $\mu_{\alpha}$, it is natural to ask whether $\alpha\mapsto\mu_{\alpha}$ is differentiable in a suitable topology. When differentiability holds, it is customary to say that the system has linear response. Such a property was first proved by Ruelle in the framework of uniformly hyperbolic systems [38] followed by Dolgopyat in the setting of partially hyperbolic systems [26]. Since then many authors studied such a phenomena and higher order differentiability in a smooth uniformly hyperbolic setting, most notably by Gouëzel and Liverani [31], Butterley and Liverani [21], among others. However, beyond uniform hyperbolicity, the situation is more subtle. For instance, even for simple piecewise smooth expanding maps if perturbations (or parameters) are transversal to the topological class, linear response can fail to hold, see the work of Baladi [11, 12], Baladi and Smania [15] and related work on unimodal maps [13, 16, 17]. On the other hand, there are other recent examples where linear response holds despite the perturbation being transversal to the topological class [9, 22]. The situation in smooth partially hyperbolic setting is not better. Despite the positive results of Dolgopyat [26], Zhang [44] studied a class of partially hyperbolic skew products where it was shown that linear response fails and that the map $\alpha\mapsto\mu_{\alpha}$ is only Hölder for certain partially hyperbolic systems.
@@
Motivated by examples of partially hyperbolic systems that do not fall within the classes considered by [26] and [44], we consider in this work skew product maps with uniformly contacting fibres. Our approach is based on a sectional transfer operator acting on families of probability measures along the fibres. The operator allows to describe invariant measures of the skew-product in terms of sample measures over the base dynamics, regardless of invertibility or non-invertibility of the base map. Hence, this approach turns out to be not only fruitful to study linear response for certain partially hyperbolic systems but also to study linear response for measures, called Bernoulli convolutions, used in fractal geometry to analyse number theoretic questions posed by Erdős [28, 40]. Under general assumptions, we establish existence and uniqueness of invariant sample measures and prove their differentiability, with respect to system parameters, in suitable topologies. In one of our applications we obtain linear response for Bernoulli convolutions [40] for all parameters $\alpha\in(0,1)$. Another application of our results yields linear response for physical measures of solenoidal attractors with intermittency, an example of a hyperbolic system which cannot be handled by traditional transfer operator techniques [19].
@@ 1.1. General setup and abstract results
As a by-product of our approach, we obtain the following linear response result for the sample measures, whose formulation is fairly standard under assumption (A3). We emphasise that the specific nature of the operator $\mathcal{K}_{\alpha}$ and of the spaces $\mathbf{E}_{\mathrm{L}}^{0},\mathbf{E}_{k}^{0}$ is immaterial for this theorem.
@@ 1.4. Applications
In this section, we apply our abstract framework to two landmark problems that serve as primary motivations for this work: the Bernoulli convolutions and the intermittent solenoid. These applications illustrate the power of our approach in handling two distinct but fundamental challenges in linear response theory. First, we provide a global differentiability result for the parameter-dependent measures of the baker map, overcoming the classical obstructions posed by singular parameters (such as Pisot numbers) that had previously restricted such results to “almost every” parameter. Second, we establish linear response for a solenoid map undergoing an intermittency transition, where the loss of uniform expansion in the base dynamics typically ruins exponential mixing and standard spectral gap arguments. By lifting the dynamics to higher-order dual spaces and utilizing the inducing schemes developed in the previous sections, we demonstrate that statistical stability can be recovered even in these notoriously “rough” parameter landscapes.
@@ 1.4.1. Bernoulli convolutions
where $(\xi_{n})_{n\geq 0}$ are i.i.d. random variables with $\mathbb{P}(a_{n}=1)=\mathbb{P}(a_{n}=-1)=\tfrac{1}{2}$, and $0<\alpha<1$. We denote by $\nu_{\alpha}$ the distribution of $Y_{\alpha}$.
A classical result of Jessen and Wintner [32] shows that $\nu_{\alpha}$ is always pure, that is, either absolutely continuous or purely singular with respect to Lebesgue measure. More precisely, for $\alpha<\tfrac{1}{2}$, the measure $\nu_{\alpha}$ is singular and supported on a Cantor set of zero Lebesgue measure. The regime $\alpha>\tfrac{1}{2}$ is considerably more subtle: Solomyak [40] proved that $\nu_{\alpha}$ is absolutely continuous for Lebesgue almost every $\alpha\in(\tfrac{1}{2},1)$, whereas Erdős [28] showed that $\nu_{\alpha}$ is singular whenever $1/\alpha$ is a Pisot number.
The connection with deterministic dynamics is provided by the baker map. Following Alexander and Yorke [6], we consider the skew-product map $T_{\alpha}\colon[-1,1]^{2}\to[-1,1]^{2}$ defined by
$T_{\alpha}(\omega,x)=(f(\omega),g_{\alpha}(\omega,x)),$
where the base map $f\colon[-1,1]\to[-1,1]$ is the doubling map
$f(\omega)=\begin{cases}2\omega+1,&\omega\in[-1,0),\\[5.69054pt] 2\omega-1,&\omega\in[0,1],\end{cases}$
@@
$g_{\alpha}(\omega,x)=\begin{cases}\alpha x-(1-\alpha),&\omega\in[-1,0),\\[5.69054pt] \alpha x+(1-\alpha),&\omega\in[0,1].\end{cases}$
This system admits a unique physical measure $\mu_{\alpha}=m\times\nu_{\alpha}$, where $m$ denotes Lebesgue measure on the base. The vertical marginal $\nu_{\alpha}$ is the unique invariant probability measure satisfying the iterated function system relation
$\nu_{\alpha}=\tfrac{1}{2}(g_{\alpha,1})_{*}\nu_{\alpha}+\tfrac{1}{2}(g_{\alpha,2})_{*}\nu_{\alpha},$
with $g_{\alpha,1}=g_{\alpha,\omega}$ for $\omega<0$ and $g_{\alpha,2}=g_{\alpha,\omega}$ for $\omega\geq 0$. Equivalently, $\nu_{\alpha}$ is the distribution of the random series
@@
where $(a_{n})_{n\geq 0}$ are i.i.d. random variables taking values in $\{-1,1\}$ with equal probability; this corresponds to a simple rescaling of the original Bernoulli convolution.
Recently, Kloeckner [35] proved that the map $\alpha\mapsto\nu_{\alpha}$ is differentiable for almost every $\alpha\in(\tfrac{1}{2},1)$ – more precisely, for a full-measure subset of the parameters considered in [40] – when tested against smooth observables, a result usually referred to as weak linear response. The coexistence of singular parameters (such as Pisot values) with regular ones has traditionally obstructed any global differentiability statement. By contrast, the framework developed in this paper allows us to obtain linear response in a strong sense, valid for all parameters $\alpha\in(0,1)$.
@@
The map $\alpha\mapsto\nu_{\alpha}$ is differentiable on $(0,1)$ with values in $C^{3}([-1,1])^{*}$.
In particular, for every $C^{3}$ observable $\phi:[-1,1]\to\mathbb{R}$, the map
$\alpha\longmapsto\int_{[-1,1]}\phi\,d\nu_{\alpha}$
is differentiable at all $\alpha\in(0,1)$. Theorem H follows as a direct application of the final item of Theorem E. Indeed, the sectional transfer operator in this setting is given by
$(\mathcal{K}_{\alpha}\bm{\nu})_{\omega}=\tfrac{1}{2}(g_{\alpha,1})_{*}\nu_{\theta_{\alpha,1}(\omega)}+\tfrac{1}{2}(g_{\alpha,2})_{*}\nu_{\theta_{\alpha,2}(\omega)}.$
Since $\mathcal{K}_{\alpha}$ preserves the closed subspace of constant sections of $\mathbf{P}$, its fixed point $\bm{\nu}_{\alpha}$ is necessarily a constant section. Consequently, the space $\mathbf{E}_{3}$ in which differentiability is obtained can naturally be identified with the dual space $C^{3}([-1,1])^{*}$. It therefore suffices to verify the assumptions of Theorem E.
(A2), (B1), (B2): These assumptions are trivially satisfied: the base map is the doubling map $f$ for all $\alpha$, with two inverse branches defined on [1,0][-1,0] and $[0,1]$, and its absolutely continuous invariant probability measure is Lebesgue measure.
(B3): Since both fibre maps are affine contractions, all the required conditions are immediately verified. The only subtlety concerns the contraction constant $\lambda$, which cannot be chosen uniformly for all $\alpha\in(0,1)$. However, as differentiability is a local property in parameter space, $\lambda$ can always be chosen locally.
(B4): This assumption is trivially verified, as there are only two weight functions, both constant and equal to $1/2$ on each inverse branch.
@@ 1.4.2. Solenoid with intermittency
Let $D\subset\mathbb{R}^{2}$ denote the unit disk and fix $\lambda\in(0,\tfrac{1}{2})$. For each $\alpha\in(0,1)$, consider the skew-product map
@@
This follows as an application of our previous results in the following way. First, note that the system satisfies the assumptions of Theorem A, which guarantees the existence of a skew-product measure as in the conclusion of that theorem, with marginal $\eta_{\alpha}$ on $\Omega$. Since $\eta_{\alpha}$ is a physical measure for the base map, it follows from [34, Theorem A] that this skew-product measure is a physical measure for $T_{\alpha}$, and thus coincides with $\mu_{\alpha}$ by uniqueness. The conclusion of Thoerem I then follows from Theorem D applied with $k=3$, together with Theorem G. Recall that the applicability of Theorem D relies on the final conclusion of Theorem G as well as on assumptions (A1)(A2) and (A4)(A5), which we now proceed to verify.
@@
It suffices to show that these quantities depend continuously on $\alpha$ in the $C^{0}$ norm. It is straightforward to verify that both the inverse branches $\theta_{\alpha,i}$ and the derivative $f_{\alpha}^{\prime}$ depend continuously on $\alpha$. Assume that $p_{\alpha,1}$ corresponds to the right-hand branch, which is uniformly expanding. It follows from (1.14) that, up to a normalising constant, $\rho_{\alpha}\circ\theta_{\alpha,1}$ coincides with $\bar{\rho}_{\alpha}\circ\theta_{\alpha,1}$. This density is bounded away from zero and infinity and depends continuously (in fact, smoothly) on the parameter $\alpha$, by [8, Theorem 3.2]. On the other hand, by the results of [43, Section 6.3], there exists a constant $C>0$ such that
@@ 2.1. Sample measures
(2.1) $W_{1}\big(({\mathcal{K}_{\alpha}}\bm{\nu})_{\omega},({\mathcal{K}_{\alpha}}\bm{\mu})_{\omega}\big)\leq\sum_{\theta\in f_{\alpha}^{-1}(\{\omega\})}\frac{\rho_{\alpha}(\theta)}{|\det Df_{\alpha}(\theta)|\rho_{\alpha}(\omega)}W_{1}((g_{\alpha,\theta})_{*}\nu_{\alpha,\theta},(g_{\alpha,\theta})_{*}\mu_{\alpha,\theta}).$
@@
$W_{1}\left((g_{\theta})_{*}\nu_{\alpha,\theta},(g_{\theta})_{*}{\mu}_{\alpha,\theta}\right)\leq\lambda\,W_{1}(\nu_{\alpha,\theta},\mu_{\alpha,\theta}),$
@@ Appendix A Dual spaces
This appendix collects the definitions and conventions for the functional spaces and norms used in the analysis. These include Lipschitz and smooth observables, their dual spaces, and the Wasserstein–$1$ distance on probability measures, together with the associated push-forward operators. We adopt the standard duality notation $\langle\nu,\phi\rangle$ for the action of a functional $\nu$ on a test function $\phi$. The norms are chosen to balance the regularity required for continuity and differentiability of transfer operators with the possible singularity of the fibre measures. All constructions are stated here for completeness and to fix notation.

### v2

@@ 1. Introduction
More precisely, given a family of dynamical systems $T_{\alpha}:M\to M$, with $\alpha$ taking values in an interval $J\subset\mathbb{R}$, it is natural to ask whether dynamical quantities are differentiable, or not, with respect to the parameter $\alpha$ in a suitable topology. In particular, if for each $\alpha$ the system $T_{\alpha}$ admits a unique physical measure $\mu_{\alpha}$, it is natural to ask whether $\alpha\mapsto\mu_{\alpha}$ is differentiable in a suitable topology. When differentiability holds, it is customary to say that the system has linear response. Such a property was first proved by Ruelle in the framework of uniformly hyperbolic systems [41] followed by Dolgopyat in the setting of partially hyperbolic systems [27]. Since then many authors studied such a phenomena and higher order differentiability in a smooth uniformly hyperbolic setting, most notably by Gouëzel and Liverani [33], Butterley and Liverani [21], among others. However, beyond uniform hyperbolicity, the situation is more subtle. For instance, even for simple piecewise smooth expanding maps if perturbations (or parameters) are transversal to the topological class, linear response can fail to hold, see the work of Baladi [11, 12], Baladi and Smania [15] and related work on unimodal maps [13, 16, 17, 24]. On the other hand, there are other recent examples where linear response holds despite the perturbation being transversal to the topological class [9, 22]. The situation in smooth partially hyperbolic setting is not better. Despite the positive results of Dolgopyat [27], Zhang [47] studied a class of partially hyperbolic skew products where it was shown that linear response fails and that the map $\alpha\mapsto\mu_{\alpha}$ is only Hölder for certain partially hyperbolic systems.
@@
Motivated by examples of partially hyperbolic systems that do not fall within the classes considered by [27] and [47], we consider in this work skew product maps with uniformly contacting fibres. Our approach is based on a sectional transfer operator acting on families of probability measures along the fibres. The operator allows to describe invariant measures of the skew-product in terms of sample measures over the base dynamics, regardless of invertibility or non-invertibility of the base map. Hence, this approach turns out to be not only fruitful to study linear response for certain partially hyperbolic systems but also to study linear response for measures, called Bernoulli convolutions, used in fractal geometry to analyse number theoretic questions posed by Erdős [29, 43]. Under general assumptions, we establish existence and uniqueness of invariant sample measures and prove their differentiability, with respect to system parameters, in suitable topologies. In one of our applications we obtain strong linear response for Bernoulli convolutions covering also the biased case addressed in [40]. Another application of our results yields linear response for physical measures of solenoidal attractors with intermittency, an example of a hyperbolic system which cannot be handled by traditional transfer operator techniques [19].
@@ 1.1. General setup and abstract results
As a by-product of our approach, we obtain the following linear response result for the sample measures. We emphasise that the specific nature of the operator $\mathcal{K}_{\alpha}$ and of the spaces $\mathbf{E}_{\mathrm{L}}^{0},\mathbf{E}_{k}^{0}$ is immaterial for this theorem.
@@ 1.4. Applications
In this section, we apply our abstract framework to two landmark problems that serve as primary motivations for this work: the Bernoulli convolutions and the intermittent solenoid. These applications illustrate the power of our approach in handling two distinct but fundamental challenges in linear response theory.
@@ 1.4.1. Bernoulli convolutions
where $(a_{n})_{n\geq 0}$ are i.i.d. random variables with $\mathbb{P}(a_{n}=1)=\mathbb{P}(a_{n}=-1)=\beta$, and $0<\alpha,\beta<1$. We denote by $\nu_{\alpha,\beta}$ the distribution of $Y_{\alpha}$.
For the classical unbiased case $\beta=1/2$, a result of Jessen and Wintner [34] shows that $\nu_{\alpha,1/2}$ is always pure, that is, either absolutely continuous or singular with respect to Lebesgue measure. More precisely, for $\alpha<1/2$, the measure $\nu_{\alpha}$ is singular and supported on a Cantor set of zero Lebesgue measure. The regime $\alpha>1/2$ is considerably more subtle: Solomyak [43] proved that $\nu_{\alpha,1/2}$ is absolutely continuous for Lebesgue almost every $\alpha\in(1/2,1)$, whereas Erdős [29] showed that $\nu_{\alpha,1/2}$ is singular whenever $1/\alpha$ is a Pisot number. The connection of this unbiased case to deterministic dynamics is provided by the baker map; see e.g. [6].
Here we address the general biased case with $\beta\in(0,1)$ studied in [40] by Peres and Solomyak. Consider the family of skew-product map $T_{\alpha,\beta}\colon[-1,1]^{2}\to[-1,1]^{2}$ defined by
$T_{\alpha,\beta}(\omega,x)=(f_{\beta}(\omega),g_{\alpha}(\omega,x)),$
where the base map $f_{\beta}\colon[-1,1]\to[-1,1]$ is given by
$f_{\beta}(\omega)=\begin{cases}\frac{1}{\beta}\omega-\frac{\beta-1}{\beta},&\omega\in[-1,2\beta-1),\\[5.69054pt] \frac{1}{1-\beta}\omega-\frac{\beta-1}{\beta},&\omega\in[2\beta-1,1];\end{cases}$
@@
$g_{\alpha}(\omega,x)=\begin{cases}\alpha x-(1-\alpha),&\omega\in[-1,2\beta-1),\\[5.69054pt] \alpha x+(1-\alpha),&\omega\in[2\beta-1,1].\end{cases}$
This system admits a unique physical measure $\mu_{\alpha,\beta}=m\times\nu_{\alpha,\beta}$, where $m$ denotes Lebesgue measure on the base. The vertical marginal $\nu_{\alpha,\beta}$ is the unique invariant probability measure of the iterated function system $\{g_{\alpha,1},g_{\alpha,2};\beta,1-\beta\}$ satisfying the relation
$\nu_{\alpha,\beta}=\tfrac{1}{\beta}(g_{\alpha,1})_{*}\nu_{\alpha,\beta}+\tfrac{1}{1-\beta}(g_{\alpha,2})_{*}\nu_{\alpha,\beta},$
where $g_{\alpha,1}=g_{\alpha}(\omega,\cdot)$ for $\omega<2\beta-1$ and $g_{\alpha,2}=g_{\alpha}(\omega,\cdot)$ for $\omega\geq 2\beta-1$. Equivalently, $\nu_{\alpha,\beta}$ is the distribution of the random series
@@
where $(a_{n})_{n\geq 0}$ are i.i.d. random variables with $\mathbb{P}(a_{n}=1)=\mathbb{P}(a_{n}=-1)=\beta$; this corresponds to a simple rescaling of the original Bernoulli convolution.
Kloeckner [37] proved that in the unbiased case the map $\alpha\mapsto\nu_{\alpha,1/2}$ is differentiable for almost every $\alpha\in(1/2,1)$ – more precisely, for a full-measure subset of the parameters considered in [43] – when tested against smooth observables, a result usually referred to as weak linear response. Recently we were made aware that in the upcoming paper [31] by Fu, the author studies weak linear response for unbiased Bernoulli convolutions against Hölder observables. In particular, it is shown that whenever the Hausdorff dimension of $\nu_{\alpha,1/2}$ is smaller than one, weak differentiability against Hölder observables fails to hold, while for $C^{1}$ observables weak linear response always hold. By contrast, the framework developed in this paper allows us to obtain linear response in a strong sense, valid for all parameters $\alpha,\beta\in(0,1)$.
@@
The probability measure $\nu_{\alpha,\beta}$ depends differentiably on $\alpha,\beta\in(0,1)$ with values in $C^{3}([-1,1])^{*}$.
Theorem H follows as a direct application of the final item of Theorem E. For that, we are going to consider the dependence on the parameters $\alpha,\beta$ separately. Nevertheless, we use the sectional transfer operator in this setting labelled by both parameters
$(\mathcal{K}_{\alpha,\beta}\bm{\nu})_{\omega}=\tfrac{1}{\beta}(g_{\alpha,1})_{*}\nu_{\theta_{\alpha,\beta,1}(\omega)}+\tfrac{1}{1-\beta}(g_{\alpha,2})_{*}\nu_{\theta_{\alpha,\beta,2}(\omega)}.$
Since $\mathcal{K}_{\alpha,\beta}$ preserves the closed subspace of constant sections of $\mathbf{P}$, its fixed point $\bm{\nu}_{\alpha,\beta}$ is necessarily a constant section. Consequently, the space $\mathbf{E}_{3}$ in which differentiability is obtained can naturally be identified with the dual space $C^{3}([-1,1])^{*}$. It therefore suffices to verify the assumptions of Theorem E.
(A2), (B1), (B2): These assumptions are trivially satisfied: the base map is $f_{\beta}$ for all $\alpha,\beta$, with two inverse branches taking values in $1,2\beta-1)$ and $2\beta-1,1]$, and its absolutely continuous invariant probability measure is Lebesgue measure.
(B3): Since both fibre maps are affine contractions, all the required conditions are immediately verified. The only subtlety concerns the contraction constant $\lambda=\alpha$, which cannot be chosen uniformly for all $\alpha\in(0,1)$. However, as differentiability is a local property in parameter space, $\lambda$ can always be chosen locally uniformly.
(B4): This assumption is trivially verified, as the weight functions are constant on each branch with values $\beta$ and $1-\beta$.
@@ 1.4.2. Solenoid with intermittency
Let $D\subset\mathbb{R}^{2}$ denote the unit disk and fix $\lambda\in(0,1/2)$. For each $\alpha\in(0,1)$, consider the skew-product map
@@
This follows as an application of our previous results in the following way. First, note that the system satisfies the assumptions of Theorem A, which guarantees the existence of a skew-product measure as in the conclusion of that theorem, with marginal $\eta_{\alpha}$ on $\Omega$. Since $\eta_{\alpha}$ is a physical measure for the base map, it follows from [36, Theorem A] that this skew-product measure is a physical measure for $T_{\alpha}$, and thus coincides with $\mu_{\alpha}$ by uniqueness. The conclusion of Thoerem I then follows from Theorem D applied with $k=3$, together with Theorem G. Recall that the applicability of Theorem D relies on the final conclusion of Theorem G as well as on assumptions (A1)(A2) and (A4)(A5), which we now proceed to verify.
@@
It suffices to show that these quantities depend continuously on $\alpha$ in the $C^{0}$ norm. It is straightforward to verify that both the inverse branches $\theta_{\alpha,i}$ and the derivative $f_{\alpha}^{\prime}$ depend continuously on $\alpha$. Assume that $p_{\alpha,1}$ corresponds to the right-hand branch, which is uniformly expanding. It follows from (1.14) that, up to a normalising constant, $\rho_{\alpha}\circ\theta_{\alpha,1}$ coincides with $\bar{\rho}_{\alpha}\circ\theta_{\alpha,1}$. This density is bounded away from zero and infinity and depends continuously (in fact, smoothly) on the parameter $\alpha$, by [8, Theorem 3.2]. On the other hand, by the results of [46, Section 6.3], there exists a constant $C>0$ such that
@@ 2.1. Sample measures
(2.1) $W_{1}\big(({\mathcal{K}_{\alpha}}\bm{\nu})_{\omega},({\mathcal{K}_{\alpha}}\bm{\mu})_{\omega}\big)\leq\sum_{\theta\in f_{\alpha}^{-1}(\{\omega\})}\frac{\rho_{\alpha}(\theta)}{|\det Df_{\alpha}(\theta)|\rho_{\alpha}(\omega)}W_{1}((g_{\alpha,\theta})_{*}\nu_{\theta},(g_{\alpha,\theta})_{*}\mu_{\theta}).$
@@
$W_{1}\left((g_{\alpha,\theta})_{*}\nu_{\theta},(g_{\alpha,\theta})_{*}{\mu}_{\theta}\right)\leq\lambda\,W_{1}(\nu_{\theta},\mu_{\theta}),$
@@ Appendix A Dual spaces
This appendix collects the definitions and conventions for the functional spaces and norms used in the analysis. These include Lipschitz and smooth observables, their dual spaces, and the Wasserstein–$1$ distance on probability measures, together with the associated push-forward operators. We adopt the standard duality notation $\langle\nu,\phi\rangle$ for the action of a functional $\nu$ on a test function $\phi$.